3)

You plan to make a simple, open topped box from a piece of sheet metal by cutting a square โ€“ of equal size โ€“ from each corner and folding up the sides

If ๐‘™ = 200๐‘š๐‘š and ๐‘ค = 150๐‘š๐‘š calculate:

a) The value of x which will give the maximum volume

b) The maximum volume of the box

c) Comment of the value obtained in part b.


4) The gain of an amplifier is found to be

๐‘ฎ = ๐Ÿ๐ŸŽ ๐ฅ๐จ๐ (๐Ÿ๐ŸŽ๐‘ฝ๐’๐’–๐’•):

The tasks are to find equations for:

a) ๐‘‘๐บ/๐‘‘๐‘‰๐‘‚๐‘ข๐‘ก

b) ๐‘‘^2๐บ/๐‘‘๐‘‰๐‘‚๐‘ข๐‘ก 2


1
Expert's answer
2021-11-01T11:45:47-0400

3.

The volume of the box can be written in the form:ย "V(x)=(L - 2\cdot x)\cdot(W- 2\cdot x)\cdot x"

Lengths and width of the box decreased that is of sheet metal byย "x"ย from each corner, and height of the box is equalย "x". We bringย "V(x)"ย to a simple form:ย "V(x)=4\cdot x^3 -2\cdot (L+W)\cdot x^2+ L\cdot W\cdot x"

To find maximum volume one compute the derivative of volume with respect toย "x"

"V^{'}_x=12\cdot x^2 - 4\cdot (L+W)\cdot x+ L\cdot W"ย and define the root of the equationย "V^{'}_x=0"ย :

"x_{1,2}=(2\cdot(L+W)\pm\sqrt{4(L+W)^2-12\cdot L\cdot W} )/12=\frac{1}{6}(L+W\pm\sqrt{L^2+W^2-L\cdot W})"

"x_1=88.38 ;\space x_2=28.29"

The first value cannot be implemented. It is clear that the box will succeed only ifย "x<W\/2"

The second value corresponds to the maximum volume shown in the figure.

Answer:ย "x=28.29 mm; max V=379037.81 mm^2"




4.

"(a) d G / d V o u t=20 * 1 * 10 /(10 * Vout * \ln 10)=20 /( Vout * \ln 10)

\\(b)
d^{2} G / d V o u t^{2}=(-20 * d(\text { Vout } * \ln 10) / d V \text { out }) /(\text { Vout } * \ln 10)^{2}
\\=-20 * 
\ln 10 /\left((\ln 10)^{2} * \text { Vout }^{2}\right)
=-20 /\left(\ln 10 * V o u t^{2}\right)"